OpenStudy (anonymous):

differentiate y=(5-/(2x^2+1))^5 i'll rewrite the question.. please.

OpenStudy (anonymous):

\[y=(5-\sqrt{2x ^{2}+1})^{5}\]

zepdrix (zepdrix):

We start with power rule, yes? \[\Large\rm y'=5\left(5-\sqrt{2x^2+1}\right)^4\color{royalblue}{\left(5-\sqrt{2x^2+1}\right)'}\]And then we apply our chain rule. (That's what the blue part is). Need to take it's derivative still.